if i feed 350 people in 150 minutes haw many people do i feed in 1 minute?
thank
you

Found 2 solutions by user_dude2008, rfer:
Answer by user_dude2008(1862)

350/150=x/1

x=7/3

x=2.3333

about 2 people

Answer by rfer(16322)

350/150=x/1
150x=350
x=2 1/3